UT 12 December 2019

OSA: D.H.Huerta

ISp: C. Bolyard

Mtn Manager. T. Forsyth

ISA: B. Rothberg (Tucson)
AO: D. Miller (Tucson)
Special Guest stars: J. Christou (Tucson), the ARCETRI players (Firenze)

Conditions are clear with scattered clouds, winds peaking up to 15 m/s

LUCI being run from RM507-5 (64 bit)

AO being run from RM507-3 (32 bit)

ARGOS takes over at night center - see google sheets for ARGOS details

ARGOS offline due to hardware issue - contacting emergency numbers, 1 finally responded and is investigating

UT 00:57 - Started up LUCIs - recovered from LUCI-2 MOS error (strain gauge)

L1 0001-0002 - 2.51 sec dark

L2 0001-0002 - 2.51 sec dark

UT 01:01 - LUCI-1 authorized only

SOUL Testing

UT 01:37 - seeing 1.7-2 - proceeding with testing, had a LUCI-1 camera wheel failure, recovered (it couldn't get to the detent)

-seeing improved for a bit (sub-arcsecond) but is rising again.

-see D. Miller's AO Wiki for details

UT 07:33 - Completed SOUL commissioning for tonight

Engineering Tasks


UT 07:15 - ARGOS is offline, hardware issue - continuing with SOUL checks, then OSCO, then Eng/D

UT 07:34 - proceeding with OSCO

-moving to HD 19332

FIrst several images bad due to not turning off sub-framing (blame Barry) - ND2 ghosting significant - counts ~15k as predicted

UT 07:50 - script seems to have timed out waiting for AOS offset -timed out after 40 seconds, now continuing

UT 08:00 - D. Miller collecting telemetry data on star for OSCO - will use with ARCETRI software for analysis

UT 08:19 - moving to OSCO #2 - HD 47587

UT 08:29 - offset failure after image L1 0157 - forcing failed state


-trying to recover without losing preset (useful for a science night)

UT 08:51 - SX SOUL in a funny state, moving off to a new AO star to continue recovery attempt

UT 08:55 - resending preset to HD 47587 - will let it run and then collect additional AO telemetry

-pupils are not being center - run AO will fail shortly

-not waiting for it to timeout

UT 09:11 - paused script - D. Miller take frames with AO routines first

sub-frame on - centered up, taking dataa

UT 09:24 - ARGOS may be back online....status shows green

-still taking some SOUL telemetry data for OSCO in the meantime

trying to move telescope manually to take an off frame - IIF and AO are in a funky state, using IE (20")

UT 09:37 - preparing to hand over to ARGOS to see if systems are fixed

-sending one last preset to clear out the funky state

-- BarryRothberg - 11 Dec 2019
Topic revision: r4 - 12 Dec 2019, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback